If a 0.3753 g sample of sodium bicarbonate (80.9%) is titrated with 0.9155N sulfuric acid, what volume of the acid should be required to produce an end point?
1
Expert's answer
2020-12-11T02:49:37-0500

Cn = n/V

n = Cn x V

n (NaHCO3) = n (H2SO4)

m (NaHCO3)=0.3753 x 0.809 = 0.3036 g

M (NaHCO3) = 84 g/mol

n (NaHCO3) = 0.3036/84.007 = 0.0036 eq

n (NaHCO3) = n (H2SO4) = 0.0036 eq

V (H2SO) = 0.9155 x 0.0036 = 0.0033 L